# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

### v1

@@ III.5 Resonance broadening
Armed with a regularised kinetic equation (Eq. 25), we may use it to compare with $N$-body simulations. This is first presented in Fig. 5, where we adjusted the considered $T_{\mathrm{reg}}$ to match the slope of the numerical simulations, near the frequency extremum (Appendix D). Interestingly, we find that Eq. (25), although heuristic, offers a satisfactory match with the numerical measurements, over the whole action domain. This highlights the fact that nonlinear broadening effects are likely responsible for the effective smoothing of the relaxation rate near the extremum of the orbital frequency.
@@
In practice, in Fig. 5, we presented a single value of ${(N,q)}$. Let us now explore the dependence of this broadening as one varies the total number of particles, $N$, and the system’s active fraction, $q$. In Fig. 7, we explore the dependence of the relaxation rate, $\mathcal{R}_{1}$, as one varies $N$ and $q$.
@@
Figure 7: Relaxation rate for profile (1), using the same convention as in Fig. 5, as one varies $q$ for fixed $N$ (top), or varies $N$ for fixed $q$ (bottom). Here, we represent only the median value over the available realisations. As one increases $N$ and decreases $q$, the relaxation rate increases, but only up to some saturation.
@@
First, in that figure, following the rescaling from Eq. (17), we find that the dependence with respect to ${(N,q)}$ is only very mild. This emphasises that, indeed, ${1/N}$ effects are responsible for this system’s relaxation. Yet, as one increases $N$ and/or lowers $q$, one can tentatively note some increase in the overall amplitude of the measured relaxation rate. However, this increase is only limited and ultimately saturates. As such, even if one keeps increasing $N$ and/or lowering $q$, the relaxation rate, $\mathcal{R}_{1}$, saturates and stops converging further toward the discontinuous prediction from the ${1/N}$ Landau equation (see Fig. 5).
@@
Let us now be more specific on this (tentative) saturation effect. For a given value of ${(N,q)}$, let us therefore estimate the system’s nonlinear timescale, $T_{\mathrm{nl}}$. This time would then be the appropriate regularisation timescale, $T_{\mathrm{reg}}$, to be considered in Eq. (25) when comparing with numerical simulations. Heuristically, the nonlinear time, $T_{\mathrm{nl}}$, corresponds to the duration during which the quasilinear assumption from Eq. (19) holds. In that view, let us therefore equate the contribution from the nonlinear term ${[\delta F,\delta H]}$, with the rate of change of the DF fluctuation, ${\partial_{t}\delta F}$, in Eq. (19). We find that ${|\delta F|/T_{\mathrm{nl}}\!\simeq\!G\,|\delta F|^{2}}$, where we assumed that ${|\delta H|\!\simeq\!G\,|\delta F|}$. Assuming that the statistics of the DF fluctuations is Poissonian (see, e.g., Eq. 41 in Chavanis, 2012c), we can take ${|\delta F|^{2}\!\simeq\!\gamma\Gamma_{\mathrm{tot}}}$, with ${\gamma\!=\!\Gamma_{\mathrm{tot}}/N}$ the vortices’ individual circulations, and ${\Gamma_{\mathrm{tot}}}$ involving the system’s active fraction. Overall, we find the scaling
@@
Interestingly, taking inspiration from Taylor and McNamara (1971); Dubin (2003), one can offer another heuristic derivation of the scaling of $T_{\mathrm{nl}}$. Rather than being driven by nonlinearities affecting the dynamics of ${\delta F}$, let us assume that it is the (finite-$N$ driven) diffusion of $F$ that drives itself the broadening of the resonance condition in Eq. (25). Dropping all factors of order unity, and keeping only the scaling dependence with respect to $N$ and $q$, one can, heuristically, rewrite Eq. (25) as
@@
Equation (28) is a self-consistent relation for the nonlinear time, $T_{\mathrm{nl}}$ (see, e.g., Taylor and McNamara, 1971; Dubin, 2003). This self-consistent relation is explored in much more detail in Appendix H. Close to the extremum of the frequency profile, we may take ${\Delta\Omega\,T_{\mathrm{nl}}\!\to\!0}$, so that Eq. (28) gives
@@
Ultimately, we recover the exact same scaling as in Eq. (26). Of course, it is no surprise that both approaches lead to the same scaling, since the kinetic equation is, by construction, sourced by nonlinear effects, i.e. one has ${\partial F/\partial t\!=\!-\langle[\delta F,\delta H]\rangle}$.
@@
We can now use numerical simulations to investigate the validity of the previous heuristic interpretations, as one varies the values of ${(N,q)}$. For a given ${(N,q)}$, we can first compute the system’s nonlinear time, $T_{\mathrm{nl}}$, following Eq. (26). For the same parameters, we can also rely on numerical simulations to measure the system’s relaxation rate, $\mathcal{R}_{1}$. Then, just like in Fig. 5, we can adjust for the value of the regularisation time, $T_{\mathrm{reg}}$ (Eq. 25), to best reproduce the slope of the relaxation rate near the extremum of the frequency profile (following Fig. 14), i.e. where broadening effects play the most important role. In Fig. 8, we explore the dependence of $T_{\mathrm{reg}}$ as a function of $T_{\mathrm{nl}}$ for a large swath of values of ${(N,q)}$.
@@
Figure 8: Dependence of the regularisation time, $T_{\mathrm{reg}}$ (Eq. 25), as a function of the nonlinear time, $T_{\mathrm{nl}}$ (Eq. 26), using numerical simulations with different values of $N$ (different colors) and $q$ (different symbols). Here, $T_{\mathrm{reg}}$ has been obtained following Appendix D to best reproduce the relaxation rate near the frequency extremum. Surprisingly, for large values of $T_{\mathrm{nl}}$, we find that $T_{\mathrm{reg}}$ saturates, i.e. the measured relaxation rate stops converging to the (sharp) kinetic prediction from Eq. (14).
@@
First, in Fig. 8, for small values of $T_{\mathrm{nl}}$, we find that increasing $T_{\mathrm{nl}}$, indeed, leads to an increase in $T_{\mathrm{reg}}$. Phrased differently, in systems with strong resonance broadening, delaying the appearance of nonlinear effects leads to a more peaked relaxation rate near the frequency extremum. Unfortunately, the recovery of the expected trend ${T_{\mathrm{reg}}\!\propto\!T_{\mathrm{nl}}}$ is only tentative here. This is due, in particular, to the reduced range in $T_{\mathrm{nl}}$ that was considered because of the numerical complexity of simulating the present system of point vortices, for a large number of pairs ${(N,q)}$ and of independent realisations (see Appendix G for details).
@@
Second, in Fig. 8, we find that for all systems with ${T_{\mathrm{nl}}\!\gtrsim\!3\!\times\!10^{5}\,T_{\mathrm{dyn}}}$, the inferred regularisation timescale, $T_{\mathrm{reg}}$, seems to saturate. Phrased differently, even if one keeps increasing the value of $N$ and/or decreasing the value of $q$, the measured relaxation rate stops converging further toward the (discontinuous) prediction from the ${1/N}$ Landau equation (Eq. 14). This is a very surprising result, and one for which we currently have no convincing explanation. Naturally, one could be worried that this saturation behaviour stems from some issues in the numerical simulations. To limit possible concerns, we performed a couple of convergence tests, as detailed in Appendix G. To summarize, we empirically find that
@@
$\frac{T_{\mathrm{reg}}}{T_{\mathrm{dyn}}}\simeq\begin{cases}\displaystyle\propto\frac{\sqrt{N}}{q}{}&\displaystyle\mathrm{if}\quad\frac{\sqrt{N}}{q}\lesssim 3\!\times\!10^{5},\\[8.61108pt] \displaystyle 3\!\times\!10^{2}{}&\displaystyle\mathrm{if}\quad\frac{\sqrt{N}}{q}\gtrsim 3\!\times\!10^{5}.\end{cases}$ (30)
@@
Of course, the various pieces of evidence presented here only offer some glimpse on the role of nonlinear effects and resonance broadening near frequency extrema, i.e. near the breakdown of the ${1/N}$ Landau equation. At this stage, we can already list three possible venues to improve upon the results presented in the present section: (i) One should go beyond the heuristic regularisation of Eq. (25) to derive a self-consistent kinetic equation in that regime; (ii) One should reproduce the effect of resonance broadening in long-range interacting systems with much cheaper computations, e.g., with the quasiperiodic cube (see, e.g., Magorrian, 2021) or with classical Heisenberg spins (see, e.g. Barré and Gupta, 2014). These would be useful to check numerically the expected scaling ${T_{\mathrm{reg}}\!\propto\!T_{\mathrm{nl}}}$; (iii) Finally, leveraging the same systems, one should explore the relaxation of systems with much larger values of $T_{\mathrm{nl}}$ to confirm the saturation effect observed in Fig. 8. All these investigations are left for future work.
@@ III.6 Long-term relaxation
Figure 9: Relaxation in $N$-body simulations of profile (1) toward the thermodynamical equilibrium (Eq. 13). Different colors correspond to different times (here the median value over the available realisations), while the dashed line is the expected Boltzmann equilibrium (see Appendix E). Relaxation is delayed by a local kinetic blocking induced by the absence of resonances in certain regions of the non-monotonic frequency profile (see Fig. 2).
In that figure, one can note that the relaxation of the system toward the central region is particularly slow. This follows from the particular shape of the present non-monotonic frequency profile (see Fig. 2). Indeed, particles within the system’s inner regions, i.e. with small actions, cannot resonantly couple to more external regions, i.e. to particles with larger actions. This was responsible for the local kinetic blocking, already visible in the ${1/N}$ Landau flux from Fig. 5. As such, these inner regions cannot relax through $1/N$ effects and must wait for the outer regions to diffuse, via ${1/N^{2}}$ effects, so that non-local resonances get populated and that the more efficient ${1/N}$ relaxation may proceed. This lack of efficient resonant couplings is responsible for the system’s particularly slow diffusion within the core regions. Phrased differently, although the system’s frequency profile is non-monotonic, there still exist regions in action space that only relax, at least initially, via ${1/N^{2}}$ effects. Such next-order contributions are the focus of the upcoming section.
@@ IV.1 Kinetic equation
We now consider profile (2), i.e. a system with a monotonic frequency profile (Fig. 2). In that case, the resonance condition in the ${1/N}$ Landau equation (Eq. 10) imposes ${J_{1}\!=\!J}$, hence leading to the vanishing of the crossed term ${\mathbf{k}\!\cdot\!\partial F_{2}/\partial\mathbf{J}}$. This is a kinetic blocking, and there is no relaxation driven by ${1/N}$ effects. [^8] Kinetic blocking at order ${1/N}$ also holds when accounting for collective effects, as described by the inhomogeneous Balescu–Lenard equation. Relaxation is greatly hampered, and occurs only through ${1/N^{2}}$ effects. Placing themselves within this regime and neglecting collective effects, Fouvry (2022) derived a closed kinetic equation describing this relaxation. The inhomogeneous ${1/N^{2}}$ Landau equation reads
@@
[…]
@@
Just like the ${1/N}$ Landau equation (Eq. 10), the ${1/N^{2}}$ Landau equation (Eq. 31) conserves the total circulation, momentum and energy (Eq. 11), satisfies an $H$-theorem for the Boltzmann entropy (Eq. 12), and vanishes exactly for the Boltzmann DF (Eq. 13).
@@
using the same notation as in Eq. (17). In that case, the ${1/N^{2}}$ Landau equation (Eq. 31) predicts that $\mathcal{R}_{2}$ should be independent of both $N$ and $q$.
@@ IV.2 Linear stability
Interestingly, in this figure, we find that the geometry of the Nyquist contours is such that, whatever the considered value of $q$, the contour never encloses the origin of the complex plane. Phrased differently, contrary to profile (1), here for profile (2), whatever the system’s active fraction, the system’s initial distribution is linearly stable. Of course, linear stability is not synonymous with weak collective amplification. Indeed, in Fig. 10, we find that for ${q\!\gtrsim\!0.004}$, the Nyquist contour gets extremely close to the origin, i.e. there exists real frequencies for which the system amplifies very significantly perturbations. In practice, in all the upcoming simulations, we limit ourselves to simulations with ${q\!\leq\!0.0007}$, so that Eq. (31) applies.

### v2

@@ III.5.3 Numerical comparison
Armed with a regularised kinetic equation (Eq. 32), we may use it to compare with $N$-body simulations. This is first presented in Fig. 5, where we adjusted the considered $T_{\mathrm{reg}}$ to match the slope of the numerical simulations, near the frequency extremum (Appendix D). Interestingly, we find that Eq. (32), although heuristic, offers a satisfactory match with the numerical measurements, over the whole action domain. This highlights the fact that nonlinear broadening effects are likely responsible for the effective smoothing of the relaxation rate near the extremum of the orbital frequency. We point out that, here, $T_{\mathrm{reg}}$ was adjusted only to match the central slope of the relaxation rate in the top panel of Fig. 5. We then used the exact same regularisation time in the bottom panel of Fig. 5 for the diffusion coefficient. There, we find that the agreement between the numerical measurement and the heuristic prediction is less convincing. Obtaining a better match requires the development of a much more involved kinetic prediction, possibly through renormalisation techniques (see, e.g., Krommes, 2002, and references therein). This will be the topic of future work.
@@
In practice, in Fig. 5, we presented a single value of (N,q){(N,q)}. Let us now explore the dependence of this broadening as one varies the total number of particles, $N$, and the system’s active fraction, $q$. In Fig. 8, we explore the dependence of the relaxation rate, $\mathcal{R}_{1}(J)$}, as one varies $N$ and $q$.
@@
Figure 8: Relaxation rate for profile (1), using the same convention as in Fig. 5, as one varies $q$ for fixed NN (top), or varies NN for fixed $q$ (bottom). Here, we represent only the median value over the available realisations. As one increases NN and decreases $q$, the relaxation rate increases, but only up to some saturation.
@@
First, in that figure, following the rescaling from Eq. (20a), we find that the dependence with respect to (N,q){(N,q)} is only very mild. This emphasises that, indeed, 1/N{1/N} effects are responsible for this system’s relaxation. Yet, as one increases $N$ and/or lowers $q$, one can tentatively note some increase in the overall amplitude of the measured relaxation rate. However, this increase is only limited and ultimately saturates. As such, even if one keeps increasing $N$ and/or lowering $q$, the relaxation rate, $\mathcal{R}_{1}(J)$}, saturates and stops converging further toward the discontinuous prediction from the 1/N{1/N} Landau equation (see Fig. 5).
III.5.4 Nonlinear timescale
@@
Let us now be more specific on this (tentative) saturation effect. For a given value of (N,q){(N,q)}, let us therefore estimate the system’s nonlinear timescale, $T_{\mathrm{nl}}$. This time would then be the appropriate regularisation timescale, $T_{\mathrm{reg}}$, to be considered in Eq. (32) when comparing with numerical simulations. Heuristically, the nonlinear time, $T_{\mathrm{nl}}$, corresponds to the duration during which the quasilinear assumption from Eq. (26) holds. In that view, let us therefore equate the contribution from the nonlinear term $\delta F,\delta H]$}, with the rate of change of the DF fluctuation, $\partial_{t}\delta F$}, in Eq. (26). We find that $|\delta F|/T_{\mathrm{nl}}\!\simeq\!G\,|\delta F|^{2}$}, where we assumed that $|\delta H|\!\simeq\!G\,|\delta F|$}. Assuming that the statistics of the DF fluctuations is Poissonian (see, e.g., Eq. (41) in Chavanis, 2012c), we can take $|\delta F|^{2}\!\simeq\!\gamma\Gamma_{\mathrm{tot}}$}, with $\gamma\!=\!\Gamma_{\mathrm{tot}}/N$} the vortices’ individual circulations, and $\Gamma_{\mathrm{tot}}$} involving the system’s active fraction. Overall, we find the scaling
@@
III.5.5 Alternative view on the nonlinear timescale
Interestingly, taking inspiration from Taylor and McNamara (1971); Dubin (2003), one can offer another heuristic derivation of the scaling of $T_{\mathrm{nl}}$. Rather than being driven by nonlinearities affecting the dynamics of $\delta F$}, let us assume that it is the (finite-$N$ driven) diffusion of $F$ that drives itself the broadening of the resonance condition in Eq. (32). This is what we explore in more detail in Appendix I recovering the regularised Landau equation from individual stochastic equations of motion. We reproduce here some of the physical insights that can be gained from this approach.
Dropping all factors of order unity, and keeping only the scaling dependence with respect to $N$ and $q$, one can, heuristically, rewrite Eq. (32) as
@@ III.5.5 Alternative view on the nonlinear timescale
Equation (37) is a self-consistent relation for the nonlinear time, $T_{\mathrm{nl}}$ (see, e.g., Taylor and McNamara, 1971; Dubin, 2003). This self-consistent relation is explored further in Appendix I.2 and the associated Fig. 19. Close to the extremum of the frequency profile, we may take $\Delta\Omega\,T_{\mathrm{nl}}\!\to\!0$}, so that Eq. (37) gives
@@
Ultimately, we recover the exact same scaling as in Eq. (35). Exact prefactors for this relation are obtained in Appendix I.1. Of course, it is no surprise that both approaches lead to the same scaling, since the kinetic equation is, by construction, sourced by nonlinear effects, i.e. one has $\partial F/\partial t\!=\!-\langle[\delta F,\delta H]\rangle$}.
III.5.6 Numerical exploration of saturation
@@
We can now use numerical simulations to investigate the validity of the previous heuristic interpretations, as one varies the values of (N,q){(N,q)}. For a given (N,q){(N,q)}, we can first compute the system’s nonlinear time, $T_{\mathrm{nl}}$, following Eq. (35). For the same parameters, we can also rely on numerical simulations to measure the system’s relaxation rate, $\mathcal{R}_{1}(J)$}. Then, just like in Fig. 5, we can adjust for the value of the regularisation time, $T_{\mathrm{reg}}$ (Eq. 32), to best reproduce the slope of the relaxation rate near the extremum of the frequency profile (following Fig. 15), i.e. where broadening effects play the most important role. In Fig. 9, we explore the dependence of $T_{\mathrm{reg}}$ as a function of $T_{\mathrm{nl}}$ for a large swath of values of (N,q){(N,q)}.
@@
Figure 9: Dependence of the regularisation time, $T_{\mathrm{reg}}$ (Eq. 32), as a function of the nonlinear time, $T_{\mathrm{nl}}$ (Eq. 35), using numerical simulations with different values of NN (different colours) and $q$ (different symbols). Here, $T_{\mathrm{reg}}$ has been obtained following Appendix D to best reproduce the relaxation rate near the frequency extremum. Surprisingly, for large values of $T_{\mathrm{nl}}$, we find that $T_{\mathrm{reg}}$ saturates, i.e. the measured relaxation rate stops converging to the (sharp) kinetic prediction from Eq. (14).
@@
First, in Fig. 9, for small values of $T_{\mathrm{nl}}$, we find that increasing $T_{\mathrm{nl}}$, indeed, leads to an increase in $T_{\mathrm{reg}}$. Phrased differently, in systems with strong resonance broadening, delaying the appearance of nonlinear effects leads to a more peaked relaxation rate near the frequency extremum. Unfortunately, the recovery of the expected trend $T_{\mathrm{reg}}\!\propto\!T_{\mathrm{nl}}$} is only tentative here. This is due, in particular, to the reduced range in $T_{\mathrm{nl}}$ that was considered because of the numerical complexity of simulating the present system of point vortices, for a large number of pairs (N,q){(N,q)} and of independent realisations (see Appendix G for details).
@@
Second, in Fig. 9, we find that for all systems with $T_{\mathrm{nl}}\!\gtrsim\!3\!\times\!10^{5}\,T_{\mathrm{dyn}}$}, the inferred regularisation timescale, $T_{\mathrm{reg}}$, seems to saturate. Phrased differently, even if one keeps increasing the value of $N$ and/or decreasing the value of $q$, the measured relaxation rate stops converging further toward the (discontinuous) prediction from the 1/N{1/N} Landau equation (Eq. 14). Following the numerical result from Fig. 9, this suggests that the 1/N{1/N} Landau equation will not be recovered in the limit $N\!\to\!+\infty$}. This is a very surprising result, and one for which we currently have no convincing explanation. Naturally, one could be worried that this saturation behaviour stems from some issues in the numerical simulations. To limit possible concerns, we performed a couple of convergence tests, as detailed in Appendix G. To summarize, we empirically find that
@@
$\frac{T_{\mathrm{reg}}}{T_{\mathrm{dyn}}}\simeq\begin{cases}\displaystyle 10^{-3}\frac{\sqrt{N}}{q}{}&\displaystyle\mathrm{if}\quad\frac{\sqrt{N}}{q}\lesssim 3\!\times\!10^{5},\\[8.61108pt] \displaystyle 3\!\times\!10^{2}{}&\displaystyle\mathrm{if}\quad\frac{\sqrt{N}}{q}\gtrsim 3\!\times\!10^{5}.\end{cases}$ (39)
@@
Of course, the various pieces of evidence presented here only offer some glimpse on the role of nonlinear effects and resonance broadening near frequency extrema, i.e. near the breakdown of the 1/N{1/N} Landau equation. At this stage, we can already list three possible venues to improve upon the results presented in the present section: (i) One should go beyond the heuristic regularisation of Eq. (32) to derive a self-consistent kinetic equation in that regime; (ii) One should reproduce the effect of resonance broadening in long-range interacting systems with much cheaper computations, e.g., with the quasiperiodic cube (see, e.g., Magorrian, 2021) or with classical Heisenberg spins (see, e.g. Barré and Gupta, 2014). These would be useful to check numerically the expected scaling $T_{\mathrm{reg}}\!\propto\!T_{\mathrm{nl}}$}; (iii) Finally, leveraging the same systems, one should explore the relaxation of systems with much larger values of $T_{\mathrm{nl}}$ to confirm the saturation effect observed in Fig. 9. All these investigations are left for future work.
@@ III.6 Long-term relaxation
Figure 10: Relaxation in $N$-body simulations of profile (1) toward the thermodynamical equilibrium (Eq. 13). Different colours correspond to different times (here the median value over the available realisations), while the dashed line is the expected Boltzmann equilibrium (see Appendix E). Relaxation is delayed by a local kinetic blocking induced by the absence of resonances in certain regions of the non-monotonic frequency profile (see Fig. 2).
In that figure, one can note that the relaxation of the system toward the central region is particularly slow. This follows from the particular shape of the present non-monotonic frequency profile (see Fig. 2). Indeed, particles within the system’s inner regions, i.e. with small actions, cannot resonantly couple to more external regions, i.e. to particles with larger actions. This was responsible for the local kinetic blocking, already visible in the 1/N{1/N} Landau flux from Fig. 5. As such, these inner regions cannot relax through $1/N$ effects and must wait for the outer regions to diffuse, via $1/N^{2}$} effects, so that non-local resonances get populated and that the more efficient 1/N{1/N} relaxation may proceed. This lack of efficient resonant couplings is responsible for the system’s particularly slow diffusion within the core regions. Phrased differently, although the system’s frequency profile is non-monotonic, there still exist regions in action space that only relax, at least initially, via $1/N^{2}$} effects. Such next-order contributions are the focus of the upcoming section.
@@ IV.1 Kinetic equation
We now consider profile (2), i.e. a system with a monotonic frequency profile (Fig. 2). In that case, the resonance condition in the 1/N{1/N} Landau equation (Eq. 10) imposes $J_{1}\!=\!J$}, hence leading to the vanishing of the crossed term $\mathbf{k}\!\cdot\!\partial F_{2}/\partial\mathbf{J}$}. This is a kinetic blocking, and there is no relaxation driven by 1/N{1/N} effects. [^8] Kinetic blocking at order 1/N{1/N} also holds when accounting for collective effects, as described by the inhomogeneous Balescu–Lenard equation. Relaxation is greatly hampered, and occurs only through $1/N^{2}$} effects. Placing themselves within this regime and neglecting collective effects, Fouvry (2022) derived a closed kinetic equation describing this relaxation. The inhomogeneous $1/N^{2}$} Landau equation reads
@@
$\times{}$ $\,\delta_{\mathrm{D}}\big(\mathbf{k}\!\cdot\!\mathbf{\Omega}\big)\,\mathbf{k}\!\cdot\!\frac{\partial}{\partial\mathbf{J}}F_{3}(\mathbf{J})\bigg],$ (40)
@@
Just like the 1/N{1/N} Landau equation (Eq. 10), the $1/N^{2}$} Landau equation (Eq. 40) conserves the total circulation, momentum and energy (Eq. 11), satisfies an $H$-theorem for the Boltzmann entropy (Eq. 12), and vanishes exactly for the Boltzmann DF (Eq. 13).
@@
using the same notation as in Eq. (20a). In that case, the $1/N^{2}$} Landau equation (Eq. 40) predicts that $\mathcal{R}_{2}(J)$} should be independent of both $N$ and $q$.
@@ IV.2 Linear stability
Interestingly, in this figure, we find that the geometry of the Nyquist contours is such that, whatever the considered value of $q$, the contour never encloses the origin of the complex plane. Phrased differently, contrary to profile (1), here for profile (2), whatever the system’s active fraction, the system’s initial distribution is linearly stable. Of course, linear stability is not synonymous with weak collective amplification. Indeed, in Fig. 11, we find that for $q\!\gtrsim\!0.004$}, the Nyquist contour gets extremely close to the origin, i.e. there exist real frequencies for which the system amplifies very significantly perturbations. In practice, in all the upcoming simulations, we limit ourselves to simulations with $q\!\leq\!0.0007$}, so that Eq. (40) applies.